OpenStudy (anonymous):

If sinθ=3/5 and q terminates in the first quadrant, find the exact value of cos2q.

jimthompson5910 (jim_thompson5910):

Hint: cos(2q) = ( cos(q) )^2 - ( sin(q) )^2

OpenStudy (anonymous):

What is q though?

jimthompson5910 (jim_thompson5910):

oh i thought it said sin(q), my bad

jimthompson5910 (jim_thompson5910):

I think q is the terminal point

OpenStudy (anonymous):

It's okai c:

OpenStudy (anonymous):

I have to figure out how much it equals so that I can get my answer :l

jimthompson5910 (jim_thompson5910):

hmm major typo because they're referring to q as an angle when they say cos(2q)

jimthompson5910 (jim_thompson5910):

does it give you a picture or diagram?

OpenStudy (anonymous):

It gives me nothing, just the possible answers

jimthompson5910 (jim_thompson5910):

ok let me have a look at that and that might help

OpenStudy (anonymous):

And all the answers are fractions

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

okai one sec

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

3/5 9/25 3/10 7/25

jimthompson5910 (jim_thompson5910):

ok if it's a typo and they mean cos(2theta), then cos(2theta) = ( cos(theta) )^2 - ( sin(theta) )^2 cos(2theta) = ( 4/5 )^2 - ( 3/5 )^2 The reason how I know that cos(theta) = 4/5 is because if you draw a right triangle with legs x and 3 with hypotenuse 5, you'll find through the pythagorean theorem that x = 4

OpenStudy (anonymous):

Oh okai, yeah I was just stuck cuhs I had no idea how I would do that :l but yeah that makes more sense, I never thought it could be a typo

jimthompson5910 (jim_thompson5910):

that's the only thing i can think of (esp if there's no pic) cause otherwise: where does q fit in?

OpenStudy (anonymous):

Haha I have no clue that's why I asked what is q? cx

jimthompson5910 (jim_thompson5910):

yeah who knows lol